Solve 2 |x+1| -2 x =|2 x -1|+1

Solve 2|x+1|-2x=|2x-1|+1

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear chilikuri

cse 1 x<-1

2|x+1|-2x=|2x-1|+1

2-(x+1)-2x=-2x+1+1

2-(x+1)=21

so x=-2

 

case 2  -1≤x<0

2|x+1|-2x=|2x-1|+1

2(x+1)-2x=-2x+1+1

x=0 

but it is not a solution because x=0 is not lie in -1≤x<0

case 3     0≤x

2|x+1|-2x=|2x-1|+1

2(x+1)-2x=2x-1+1

 2(x+1)=2(x+1)

which is always true so x≥0 is its solution


Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin